# Gaining More Insight into Neural Semantic Parsing with Challenging Benchmarks

### v1

@@ 1.   Introduction
In the current release of the PMB, the data splits were randomly decided and considered "standard". However, this random split may result in overlap and imprecise error estimates (Søgaard et al., 2021) and and cannot adequately represent the distribution of the dataset. For instance, the sentence “I like chocolate ice cream!” is allocated to the training set, while the very similar sentence “I like chocolate ice cream.” is assigned to the test set. Equally alarmingly, some instances in the development and test sets mirror those in the training set, potentially skewing parser evaluations. Consequently, this may lead to parser evaluation results that are overly optimistic. We completely agree with Groschwitz et al. (2023), whose paper’s title "AMR Parsing is far from solved" hits the nail on the head, and even goes beyond Abstract Meaning Representation (AMR) and also applies to DRS. We think the current PMB test set lacks difficulty, because it puts emphasis on brief and simplistic sentences with an average length of less than ten words. The reason for this is that all instances of the test set have the “gold” annotation status, obtained via intensive manual correction, and the longer a document the harder it is to get an error-free annotation for it.
@@ 3.   Improving Semantic Evaluation
In this section we outline the methods to create better test sets. Besides the standard test set created with a different data split, we also show how we built additional challenge test sets.
@@ 3.1.   Splitting Data Systematically
Table 1: Distribution of train, development, and test sets in the PMB using the systematic split, together with two challenge sets. The average sentence length of each set are provided in brackets.
@@ 4.3.1.   Standard Test
Figure 3: Distribution of word overlap rates between train and test sets in EN, DE, NL, IT. Lower overlap rates signify fewer words occurring in both train and test sets. The overlap of train and dev sets is shown in Appendix A.2.
@@ 4.3.3.   Challenge Test Sets
The results of the models on the challenge test sets are shown in Tables 8 and 9. The performance on the long-text test set is significantly inferior, marked by a high incidence of ill-formed outputs. the most pronounced drop is observed in ByT5, which shows a reduction of 86% compared to the standard test set. In the generation task, although truncation does not hugely impact on evaluation, the models still grapple with long sequences, reflecting decreases of at least 29.9, 11.9, and 16.2 across three metrics. Notably, neural models struggle with the long set, primarily because their tokenization significantly amplifies both input and output lengths. For example, while the average sentence lengths in the long set stand at 61 for text and 253 for DRS, these numebrs increase to 98 and 503 after BPE tokenization (mT5, mBART, and DRS-MLM) and even further to 410 and 1370 with character-level tokenization (ByT5). Obviously, these models can not handle such long sequences as effectively as the short sequences in the standard test.
@@ Appendix A.1 Pseudo-code for CCG recombination
Appendix A.2 Distribution of word overlap
The word overlap distribution between train and development sets is consistent with train and test, as shown in Figure 4.

### v2

@@ 1.   Introduction
In the current release of the PMB, the data splits were randomly decided and considered "standard". However, this random split may result in overlap and imprecise error estimates (Søgaard et al., 2021) and and cannot adequately represent the distribution of the dataset. For instance, the sentence “I like chocolate ice cream!” is allocated to the training set, while the very similar sentence “I like chocolate ice cream.” is assigned to the test set. Equally alarmingly, some instances in the development and test sets mirror those in the training set, potentially skewing parser evaluations. Consequently, this may lead to parser evaluation results that are overly optimistic. We completely agree with opitz-frank-2022-better and Groschwitz et al. (2023), who both argue that "AMR Parsing is far from solved" hits the nail on the head, and even goes beyond Abstract Meaning Representation (AMR) and also applies to DRS. We think the current PMB test set lacks difficulty, because it puts emphasis on brief and simplistic sentences with an average length of less than ten words. The reason for this is that all instances of the test set have the “gold” annotation status, obtained via intensive manual correction, and the longer a document the harder it is to get an error-free annotation for it.
@@ 3.   Improving Semantic Evaluation
In this section we outline the methods to create better test sets. Besides the standard test set created with a different data split, we also show how we built additional challenge test sets. The resulting data set will be released as PMB 5.0.0111The release is available at https://pmb.let.rug.nl/releases/.
@@ 3.1.   Splitting Data Systematically
Table 1: Distribution of train, development, and test sets in PMB 5.0.0 using the systematic split, together with two challenge sets. The average sentence length of each set are provided in brackets.
@@ 4.3.1.   Standard Test
Figure 3: Distribution of word overlap rates between train and test sets in EN, DE, NL, IT. Lower overlap rates signify fewer words occurring in both train and test sets.
@@ 4.3.3.   Challenge Test Sets
The results of the models on the challenge test sets are shown in Tables 8 and 9. The performance on the long-text test set is significantly inferior, marked by a high incidence of ill-formed outputs [^3] SMATCH employs a hill-climbing technique to identify the optimal match, which may introduce inaccuracies when evaluating the output of the model for long texts opitz-frank-2022-better. In this case, the results for long texts should be considered as reference only.. The most pronounced drop is observed in ByT5, which shows a reduction of 86% compared to the standard test set. In the generation task, although truncation does not hugely impact on evaluation, the models still grapple with long sequences, reflecting decreases of at least 29.9, 11.9, and 16.2 across three metrics. Notably, neural models struggle with the long set, primarily because their tokenization significantly amplifies both input and output lengths. For example, while the average sentence lengths in the long set stand at 61 for text and 253 for DRS, these numebrs increase to 98 and 503 after BPE tokenization (mT5, mBART, and DRS-MLM) and even further to 410 and 1370 with character-level tokenization (ByT5). Obviously, these models can not handle such long sequences as effectively as the short sequences in the standard test.